# Multi-granularity Interest Retrieval and Refinement Network for Long-Term User Behavior Modeling in CTR Prediction

### v1

@@ Abstract.
To this end, we propose Multi-granularity Interest Retrieval and Refinement Network (MIRRN). Specifically, we first construct queries based on behaviors observed at different time scales to obtain subsequences, each capturing users’ interest at various granularities. We then introduce an noval multi-head Fourier transformer to efficiently learn sequential and interactive information within the subsequences, leading to more accurate modeling of user interests. Finally, we employ multi-head target attention to adaptively assess the impact of these multi-granularity interests on the target item. Extensive experiments have demonstrated that MIRRN significantly outperforms state-of-the-art baselines. Furthermore, an A/B test shows that MIRRN increases the average number of listening songs by 1.32% and the average time of listening songs by 0.55% on a popular music streaming app. The implementation code is publicly available at https://github.com/psycho-demon/MIRRN.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. INTRODUCTION
Figure 1. The left side of the picture illustrates a user’s behavior sequence, with black sneakers as the target item. The middle column depicts the user’s multi-granularity interests. The right column displays the subsequences related to the user’s multi-granularity interests.
[…]
To address the aforementioned challenges, we propose Multi-granularity Interest Retrieval and Refinement Network (MIRRN). Firstly, we propose multi-granularity interest retrieval module (MIRM) to capture diverse user interests. MIRM constructs queries based on the behaviors observed at different time scales, which contain the target, local, and global information of the behavior sequence, respectively. Using these queries, MIRM adopts a fast search method based on SimHash to handle behavior sequences of thousands of lengths, obtaining subsequences that each contains interest of different granularities. Secondly, we propose Behavior Sequence Refinement Module (BSRM) to extract relational information in subsequences. To incorporate sequential information and efficiently mine interactive information, BSRM employs Target-Aware Position Encoding and Multi-Head Fourier Transformer (MHFT), respectively. MHFT utilizes the fast Fourier transform (FFT) to replace the complex convolution operation in the time domain, thereby enabling interaction with a low time complexity of $O(nlogn)$. Additionally, MHFT introduces matrix multiplication and multi-head mechanism to enhance interaction and focus on information in different subspaces, while greatly reducing the number of parameters. Finally, we propose multi-granularity interest activation module (MIAM) using the multi-head target attention to adaptively learn the influence of different granular interests on the target item, obtaining the final user representation.
@@
  • In our proposed MIRRN, MIRM constructs various queries and then obtains subsequences through a fast retrieval method. This approach enables MIRM to effectively capture multi-granularity interests. To the best of our knowledge, we are the first to successfully model the user’s multi-granularity interests in long-term user behavior modeling problem.
@@
  • Extensive experiments have demonstrated that MIRRN achieves better performance compared to state-of-the-art baselines. We verify the validity of each module through comprehensive ablation studies. Furthermore, an online A/B testing conducted on a popular music streaming app. further reinforces the effectiveness of our MIRRN in real industrial scenarios.
@@ 2.1. Click-Through-Rate Prediction
Early CTR models primarily concentrate on exploiting feature interactions. Pioneering studies are relatively shallow such as factorization machines(FM) (Rendle, 2010) and its variants (Pan et al., 2016; Juan et al., 2016; Pan et al., 2018; Cheng et al., 2014). With the widespread success of deep learning, various studies such as Wide&Deep (Cheng et al., 2016), PNN (Qu et al., 2016), DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017, 2021c), successfully combine deep and shallow models. Moreover, several models (Xiao et al., 2017; He and Chua, 2017; Lian et al., 2018; Yu et al., 2023, 2024) have explored more complex neural networks, such as xDeepFM (Lian et al., 2018) further introduces compressed interaction networks to explicitly model high-order feature interactions.
@@ 3.2.2. Local-Aware Search Unit (LASU)
Specifically, we represent the user’s recent behavior sequence as $S_{r}=[e_{L-J},\dots,e_{L}]$, where $J$ is the length of recent behavior sequence. Using average pooling to construct queries is a good choice. However, considering that $J$ is very small, we can further model $S_{r}$ to enhance the representation ability of the query with slightly increasing the computational cost. Considering that each behavior is a carrier of current potential interest, we use a GRU to extract a series of hidden states $H_{r}\in\mathbb{R}^{J\times d}$ from $S_{r}$. We then combine $H_{r}$ and $S_{r}$ to get the query $q_{l}$. $Avg()$ is the average pooling.
@@ 3.2.3. Global-Aware Search unit (GASU)
In this unit, we expand the focus to encompass the entire behavior sequence to consider the user’s more global and stable interest. Constructing a query using the user’s entire long-term behavior sequence is challenging. On the one hand, directly average pooling the long sequence may result in poor query representation due to its length and the noise within the sequence. On the other hand, constructing the query with more complex modeling of the long sequence cannot meet the strict time limits of an online system. Additionally, for new users with relatively short behavior sequences, it is difficult to obtain their stable global and stable preferences.
To address the above issues, we employ clustering methods to aggregate the item set $M$ into several clusters using pre-trained item embeddings. In this search unit, we construct a query by mapping the target item to the nearest cluster centroid. This query can efficiently and roughly capture the user’s global preferences. Additionally, leveraging the semantic information of pre-trained embeddings helps to partially solve the user cold start problem.
@@ 3.3.1. Target-Aware Position Encoding (TAPE)
Previous work has rarely emphasized the order of behaviors in SBSs, possibly sorting them by relevance in the retrieval module. However, we believe that the sequential information within SBSs plays a crucial role, much like the original sequence, in helping us better understand and learn the user’s interests. Therefore, we preserve the original sequence of behaviors in SBS and use position encoding to supplement the sequential relationship, compensate for the information loss of the retrieval module, and enhance the expressive power of $S_{*}$. We use the relative position of each behavior in $S_{*}$ to the target item to construct a position encoding, termed target-aware position encoding. Rather than the widely used chronological order encoding (Kang and McAuley, 2018), this method better emphasizes the attention on the target information. Specifically, we propose the Target-Aware Position Encoding $TPE\in\mathbb{R}^{L\times d}$. We then convert the retrieved indices $I$ in equation (6) to target-aware indices $I^{\prime}$. Finally, we obtain the target-aware position embeddings through a look-up process and incorporate them into $S_{*}$:
(15) $S_{*}=S_{*}+lookup(TPE,I^{\prime})).$
@@ 3.3.2. Multi-Head Fourier Transformer (MHFT)
After applying the above TAPE, our model has learned the sequential information in SBSs. To efficiently learn the interactive information with less additional computational cost, we leverage the efficient Fourier Transformer. According to the convolution theorem (Soliman and Srinath, 1990), the product operator in the frequency domain is equivalent to global circular convolution in the time domain. This implies that complex convolution operations can be implemented through fast Fourier transform (FFT) with a time complexity of $O(KlogK)$ and multiplication operations (Lee-Thorp et al., 2021; Rao et al., 2021). Compared with the self-attention mechanism with a time complexity of $O(K^{2})$, this method explores interactive information while improving efficiency.
Specifically, we perform Fourier transform along the sequence length dimension to transform $S_{*}$ to $X\in\mathbb{C}^{K\times d}$ which represents the frequency domain representation of $S_{*}$. We then implement the convolution fusion of behaviors and mine interactive information by performing a dot product with the complex parameter matrix $W\in\mathbb{C}^{K\times d}$ (Zhou et al., 2022). Finally, we obtain the full convolution sequence representation $\widetilde{S_{*}}\in\mathbb{R}^{K\times d}$ through inverse fast Fourier transform (IFFT). The formulas are as follows:
@@
Additionally, inspired by the multi-head self attention mechanism, we further propose the Multi-Head Fourier Transformer. To achieve better information fusion, we first use matrix multiplication instead of dot product operation in equation (16), using $W\in\mathbb{C}^{K\times d\times d}$ instead of $W\in\mathbb{C}^{K\times d}$. Then we shares weights $W\in\mathbb{C}^{d\times d}$ among different tokens, to adaptively handle sequences of different lengths while reducing the parameter count. To mine the interactive information from different subspaces, we divide $X\in\mathbb{C}^{K\times d}$ into $n$ parts $x_{i}\in\mathbb{C}^{K\times d/n}$ and diagonalize the weight $W\in\mathbb{C}^{d\times d}$ into $n$ blocks $W_{i}\in\mathbb{C}^{d/n\times d/n}$. Each block of $X$ and the diagonal matrix is equivalent to a head of the multi-head self attention. The diagonal weights can be computed in parallel. We implement them through an MLP with an activation function $\sigma$ to introduce nonlinear relationships. The formulas are as follows:
@@ 3.6.2. Behavior Sequence Refinement Module (BSRM)
Table 1. Time complexity and parameter count for MHSA, FM and MHFT.
@@ 4.3. Ablation Study
(3)We further verify the effectiveness of two key components in BSRM, TAPE and MHFT, which are responsible for mining sequential and interactive information respectively. The results in Table 5 demonstrate that both the sequential information in TAPE and the interactive information in MHFT positively influence performance.
@@ 4.4. Inference Time Analysis
To demonstrate the practical value of our model, we compare the inference time across various models, including DIN, DIEN, SIM, ETA, SDIM, and MIRRN. Notably, DIN and DIEN utilize the most recent 100 behaviors, whereas the other models leverage the most recent 300 behaviors. As shown in Figure 3, SIM and SDIM exhibit relatively low inference times among the long sequence modeling methods. Our method introduces additional modules, leading to a slight increase in inference time. However, compared to the fastest baseline, our model only increases inference time by 12.0%, 11.9%, and 10.1% across the three datasets, while significantly improving AUC by 1.45%, 1.51%, and 0.58% over the strongest baseline.
@@ 4.5. Online A/B Test
We further conduct an online A/B test on the core playlist recommendation channel of on a popular music streaming app, which is utilized by millions of users daily. Figure 4 gives a simple illustration of the deployed system, which encompasses both the offline training and online inference processes. During the offline phase, the system logs hundreds of millions of user interactions to refine the prediction model, subsequently uploading the updated model to the online server for real-time user inference. In the online phase, upon receiving a new request, the online servers retrieve the relevant features and forward them to the ranking server, which then generates and presents a personalized recommendation list to users.
Figure 4. A simple illustration of the deployed playlist recommendation system on a popular music streaming app.
@@ 5. CONCLUSION
In this paper, we proposed the Multi- granularity Interest Retrieval and Refinement Network (MIRRN) for long-term user behavior modeling in CTR prediction. To the best of our knowledge, MIRRN is the first to successfully capture the user’s multi-granularity interests. Specifically, we empolyed Multi-granularity Interest Retrieval Module to extract subsequences containing interests of different granularities. We then utilized Target-Aware Position Encoding and Multi-Head Fourier Transformer to efficiently and effectively mine sequential information and interactive information within these subsequences. Finally, we used multi-head target attention to adaptively learn the influence of different granular interests on the target item. Experiments on three public datasets, along with an online A/B test validated the effectiveness of MIRRN. In the future, we plan to consider user interests at additional granularities tailored to different scenarios and explore more efficient retrieval methods to further reduce time complexity.

### v2

@@ Abstract.
To this end, we propose Multi-granularity Interest Retrieval and Refinement Network (MIRRN). Specifically, we first construct queries based on behaviors observed at different time scales to obtain subsequences, each capturing users’ interest at various granularities. We then introduce an noval multi-head Fourier transformer to efficiently learn sequential and interactive information within the subsequences, leading to more accurate modeling of user interests. Finally, we employ multi-head target attention to adaptively assess the impact of these multi-granularity interests on the target item. Extensive experiments have demonstrated that MIRRN significantly outperforms state-of-the-art baselines. Furthermore, an A/B test shows that MIRRN increases the average number of listening songs by 1.32% and the average time of listening songs by 0.55% on the Huawei Music App. The implementation code is publicly available at https://github.com/USTC-StarTeam/MIRRN.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709438isbn: 979-8-4007-1245-6/25/08ccs: Information systemsccs: Information systems Recommender systems
@@ 1. INTRODUCTION
Figure 1. The left side of the picture illustrates a user’s behavior sequence, with black sneakers as the target item. The middle column illustrates the microcosm of users’ interests at different granularities, which represents the multi-granularity queries we aim to construct. The right column displays the subsequences related to the multi-granularity interests.
[…]
@@
To address the aforementioned challenges, we propose Multi-granularity Interest Retrieval and Refinement Network (MIRRN). Firstly, we propose Multi-granularity Interest Retrieval Module (MIRM) to capture diverse user interests. MIRM constructs queries based on the behaviors observed at different time scales, which contain the target, local, and global information of the behavior sequence, respectively. Using these queries, MIRM adopts a fast search method based on SimHash to handle behavior sequences of thousands of lengths, obtaining subsequences that each contains interest of different granularities. Secondly, we propose Behavior Sequence Refinement Module (BSRM) to extract relational information in subsequences. To incorporate sequential information and efficiently mine interactive information, BSRM employs Target-Aware Position Encoding (TAPE) and Multi-Head Fourier Transformer (MHFT), respectively. MHFT utilizes the fast Fourier transform (FFT) to replace the complex convolution operation in the time domain, thereby enabling interaction with a low time complexity of $O(nlogn)$. Additionally, MHFT introduces matrix multiplication and multi-head mechanism to enhance interaction and focus on information in different subspaces, while greatly reducing the number of parameters. Finally, we propose Multi-granularity Interest Activation Module (MIAM) using the multi-head target attention to adaptively learn the influence of different granular interests on the target item, obtaining the final user representation.
@@
  • In our proposed MIRRN, MIRM construct diverse queries based on behaviors observed at different time scales to obtain subsequences. This approach enables MIRM to effectively capture multi-granularity interests. To the best of our knowledge, we are the first to successfully capture, refine, and integrate user’s multi-granularity interests in long-term user behavior modeling.
@@
  • Extensive experiments have demonstrated that MIRRN achieves better performance compared to state-of-the-art baselines. We verify the validity of each module through comprehensive ablation studies. Furthermore, an online A/B testing conducted on the Huawei Music App further reinforces the effectiveness of MIRRN in real industrial scenarios.
@@ 2.1. Click-Through-Rate Prediction
Early CTR models primarily concentrate on exploiting feature interactions. Pioneering studies are relatively shallow such as factorization machines(FM) (Rendle, 2010) and its variants (Pan et al., 2016; Juan et al., 2016; Pan et al., 2018). With the great success of deep learning, various studies such as Wide&Deep (Cheng et al., 2016), PNN (Qu et al., 2016), DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017, 2021c), successfully combine deep and shallow models. Moreover, several models (Xiao et al., 2017; He and Chua, 2017; Lian et al., 2018; Yu et al., 2023, 2024) have explored more complex neural networks, such as xDeepFM (Lian et al., 2018) further introduces compressed interaction networks to explicitly model high-order feature interactions.
@@ 3.2.2. Local-Aware Search Unit (LASU)
Specifically, we represent the user’s recent behavior sequence as $S_{r}=[e_{L-J},\dots,e_{L}]$, where $J$ is the length of recent behavior sequence. Using average pooling to construct queries is a good choice. Additionally, considering that $J$ is very small, we can further model $S_{r}$ to enhance the representation ability of the query with slightly increasing the computational cost. Considering that each behavior is a carrier of current potential interest, we use a GRU to extract a series of hidden states $H_{r}\in\mathbb{R}^{J\times d}$ from $S_{r}$. We then combine $H_{r}$ and $S_{r}$ to get the query $q_{l}$. $Avg()$ is the average pooling.
@@ 3.2.3. Global-Aware Search unit (GASU)
In this unit, we expand the focus to encompass the entire behavior sequence to consider the user’s more global and stable interest. Constructing a query using the user’s entire long-term behavior sequence is challenging. On the one hand, directly average pooling the long sequence may result in poor query representation due to its length and the noise within the sequence. On the other hand, constructing the query with more complex modeling of the long sequence cannot meet the strict time limits of an online system. Additionally, for new users with relatively short behavior sequences, it is difficult to obtain their global and stable preferences.
To address the above issues, we employ clustering methods to aggregate the item set $M$ into several clusters using pre-trained item embeddings. Then we map the target item to the nearest cluster centroid in order to construct a query, which effectively and approximately captures the global preferences. Additionally, leveraging the semantic information of pre-trained embeddings helps to partially solve the user cold start problem.
@@ 3.3.1. Target-Aware Position Encoding (TAPE)
Previous work has rarely emphasized the order of behaviors in SBSs, possibly sorting them by relevance in the retrieval module. However, we believe that the sequential information within SBSs plays a crucial role, much like the original sequence, in helping us better understand and learn the user’s interests. Therefore, we preserve the original sequence of behaviors in SBS and use position encoding to supplement the sequential relationship, compensate for the information loss of the retrieval module, and enhance the expressive power of $S_{*}$. We use the relative position of each behavior in $S_{*}$ to the target item to construct the position encoding. Rather than the widely used chronological order encoding (Kang and McAuley, 2018), this method better emphasizes the attention on the target information. Specifically, we propose the Target-Aware Position Encoding $TAPE\in\mathbb{R}^{L\times d}$. We then convert the retrieved indices $I$ in equation (6) to target-aware indices $I^{\prime}$. Finally, we obtain the target-aware position embeddings through a look-up process and incorporate them into $S_{*}$:
(15) $S_{*}=S_{*}+lookup(TAPE,I^{\prime})).$
@@ 3.3.2. Multi-Head Fourier Transformer (MHFT)
After applying the above TAPE, our model has obtained the sequential information in SBSs. To efficiently learn the interactive information with less additional computational cost, we leverage the efficient Fourier Transformer. According to the convolution theorem (Soliman and Srinath, 1990), the product operator in the frequency domain is equivalent to global circular convolution in the time domain, which is able to extract interactive information and refine the sequence representation. It also implies that complex convolution operations can be implemented through fast Fourier transform (FFT) with a time complexity of $O(KlogK)$ and multiplication operations (Lee-Thorp et al., 2021; Rao et al., 2021). Compared with the self-attention mechanism with a time complexity of $O(K^{2})$, this method explores interactive information while improving efficiency.
Specifically, we perform Fourier transform along the sequence length dimension to transform $S_{*}$ to $X\in\mathbb{C}^{K\times d}$ which represents the frequency domain representation of $S_{*}$. We then implement the convolution fusion of behaviors and mine interactive information by performing the element-wise multiplication with a complex parameter matrix $W\in\mathbb{C}^{K\times d}$ (Zhou et al., 2022). Finally, we obtain the full convolution sequence representation $\widetilde{S_{*}}\in\mathbb{R}^{K\times d}$ through inverse fast Fourier transform (IFFT). The formulas are as follows:
@@
Additionally, inspired by the multi-head self attention mechanism, we further propose the Multi-Head Fourier Transformer. To achieve better information fusion, we first use matrix multiplication instead of element-wise multiplication in equation (16), using $W\in\mathbb{C}^{K\times d\times d}$ instead of $W\in\mathbb{C}^{K\times d}$. Then we shares weights $W\in\mathbb{C}^{d\times d}$ among different tokens, to adaptively handle sequences of different lengths while reducing the parameter count. To mine the interactive information from different subspaces, we divide $X\in\mathbb{C}^{K\times d}$ into $n$ parts $x_{i}\in\mathbb{C}^{K\times d/n}$ and diagonalize the weight $W\in\mathbb{C}^{d\times d}$ into $n$ blocks $W_{i}\in\mathbb{C}^{d/n\times d/n}$. Each block of $X$ and the diagonal matrix is equivalent to a head of the multi-head self attention. The diagonal weights can be computed in parallel. We implement them through an MLP with an activation function $\sigma$ to introduce nonlinear relationships. The formulas are as follows:
@@ 3.6.2. Behavior Sequence Refinement Module (BSRM)
Table 1. Time complexity and parameter count for MHSA, FT and MHFT.
@@ 4.3. Ablation Study
(3)We further verify the effectiveness of two key components, TAPE and MHFT, in BSRM, which are responsible for mining sequential and interactive information respectively. The results in Table 5 demonstrate that both sequential information in TAPE and interactive information in MHFT positively influence performance.
@@ 4.4. Inference Time Analysis
To demonstrate the practical value of our model, we compare the inference time across various models, including DIN, DIEN, SIM, ETA, SDIM, and MIRRN. As shown in Figure 3, SIM and SDIM exhibit relatively low inference times among the long sequence modeling methods. Our method introduces additional modules, leading to a slight increase in inference time. However, compared to the fastest baseline, our model only increases inference time by 12.0%, 11.9%, and 10.1% across the three datasets, while significantly improving AUC by 1.45%, 1.51%, and 0.58% over the strongest baseline.
@@ 4.5. Online A/B Test
We further conduct an online A/B test on the core playlist recommendation channel of the Huawei Music App, which is utilized by millions of users daily. Figure 4 gives a simple illustration of the deployed system, which encompasses both the offline training and online inference processes. During the offline phase, the system logs hundreds of millions of user interactions to refine the prediction model, subsequently uploading the updated model to the online server for real-time user inference. In the online phase, upon receiving a new request, the online servers retrieve the relevant features and forward them to the ranking server, which then generates and presents a personalized recommendation list to users.
Figure 4. A simple illustration of the deployed playlist recommendation system in Huawei.
@@ 5. CONCLUSION
In this paper, we proposed the Multi-granularity Interest Retrieval and Refinement Network (MIRRN) for long-term user behavior modeling in CTR prediction. To the best of our knowledge, MIRRN is the first to successfully capture, refine, and integrate the user’s multi-granularity interests. Specifically, we constructed multi-granularity queries, extracting subsequences that capture diverse interests. We then utilized Target-Aware Position Encoding and Multi-Head Fourier Transformer to efficiently and effectively mine sequential information and interactive information within these subsequences, obtaining the refined interest vectors. Finally, we employed multi-head target attention to adaptively learn the influence of different granular interests on the target item. Experiments on three public datasets, along with an online A/B test validated the effectiveness of MIRRN. In the future, we plan to consider user interests at additional granularities tailored to different scenarios and explore more efficient retrieval methods to further reduce time complexity.
6. ACKNOWLEDGMENTS
This work is supported by the National Natural Science Foundation of China (No. U23A20319, 62472394, and 62202443). And we thank MindSpore for the partial support of this work, which is a new deep learning computing framework.